# Agnostic Process Tomography

### v1

@@ 5.7 Proper APT
where $x\in\mathbb{R}^{4D^{2}}$. Moreover, we can compute the normal cone of the polyhedral set $C_{2}=\{x\in\mathbb{R}^{4D^{2}}:\tilde{A}x=\tilde{b}\}$:
@@
=$=\{g:g^{\intercal}(x-y)\geq 0,\quad\forall y\;\text{ s.t. }\tilde{A}y=\tilde{b}\}$ (5.221)
=$=\{g:g^{\intercal}z\geq 0,\quad\forall z\in\mathrm{ker}(\tilde{A})\}$ (5.222)
$=(\mathrm{ker}(\tilde{A}))^{\perp}$ (5.223)
$=\mathrm{im}(\tilde{A}^{\intercal}).$ (5.224)
Here, the second line is the definition of the normal cone. The third line follows because $x-y$ is in the kernel of the matrix $\tilde{A}$ since both $x,y\in C_{2}$. The fifth line follows by the standard result in linear algebra that the orthogonal complement of the kernel space is the image of the transpose.
@@
$\frac{\|\hat{\Psi}_{1}-\hat{\Psi}_{2}\|_{F}}{2}\leq\frac{\|\hat{\Psi}_{1}-\hat{\Psi}\|_{F}}{2}+\epsilon_{2},\quad\forall\text{ CPTP maps }\Psi.$ (5.228)
@@
Even though $\hat{\Psi}_{2}$ has desirable distance properties, as it does not achieve exact convergence to the projection, it may not correspond to a CPTP map. However, as our final projection was onto the set $C_{2}$, $\hat{\Psi}_{2}$ is Hermitian and satisfies the affine constraints. Thus, the only remaining condition necessary to obtain a proper hypothesis is to obtain a positive semidefinite hypothesis. To ensure that our final hypothesis is PSD while still retaining the other properties, we take inspiration from [168] and mix it with the maximally mixed state. In particular, for a suitably chosen $p>0$, our output is
@@
where $I$ is the $D\times D$ identity operator. To understand the intuition behind mixing, we first note that the maximally mixed state can be viewed as the Fourier coefficient matrix of a CPTP map that applies a Pauli operator uniformly at random. Thus, it must satisfy all required constraints. Then, we note that if $p$ is taken large enough, $\hat{\Psi}_{3}$ also satisfies the constraints. This can be seen by noting that the convex combination of two points satisfying the affine constraints will also satisfy them. For the positive semidefiniteness, we note that $\hat{\Psi}_{2}$ is Hermitian, and we only need to take $p$ large enough to make the most negative eigenvalue of $\hat{\Psi}_{2}$ map to 00 in the eigenspectrum of $\hat{\Psi_{3}}$. Using the fact that $\hat{\Psi}_{2}$ is not too far from a PSD matrix (Equation 5.226), we will then show that $p$ does not need to be too large.
@@
We can now upper bound $|\lambda^{\star}|$ as follows. First, recall that projection of a Hermitian operator onto the positive semidefinite cone is well-known (see, e.g., [147]). In particular, this projection works by taking the positive component of the Hermitian operator’s eigenspectum and setting the negative eigenvalues to 00. For a Hermitian $A$, we denote this by $(A)^{+}$. Then, we have
@@
$=\frac{p}{2}\norm{\hat{\Psi}_{2}-\frac{I}{D}}_{F}$ (5.238)
$\leq\frac{p}{2}\left(\|\hat{\Psi}_{2}\|_{F}+\norm{\frac{I}{D}}_{F}\right)$ (5.239)
$\leq\frac{p}{2}\left(\|\hat{\Psi}_{2}-\hat{\Psi}^{*}\|_{F}+\|\hat{\Psi}^{*}\|_{F}+\norm{\frac{I}{D}}_{F}\right)$ (5.240)

### v2

@@ 5.7 Proper APT
where $x\in\mathbb{R}^{4D^{2}}$. Moreover, we can compute the normal cone of the polyhedral set $C_{2}=\{x\in\mathbb{R}^{4D^{2}}\mathrel{\mathop{\ordinarycolon}}\tilde{A}x=\tilde{b}\}$:
@@
$=\{g\mathrel{\mathop{\ordinarycolon}}g^{\intercal}(x-y)\geq 0,\qquad\forall y\;\text{ s.t. }\tilde{A}y=\tilde{b}\}$ (5.221)
[…]
$)\displaystyle=(\mathrm{ker}(\tilde{A}$$)^{\perp}$ (5.223)
[…]
Here, the second line is the definition of the normal cone. The third line follows because xyx-y is in the kernel of the matrix $\tilde{A}$ since both $x,y\in C_{2}$. The fifth line follows by the standard result in linear algebra that the orthogonal complement of the kernel space is the image of the transpose.
@@
$\frac{\|\hat{\Psi}_{1}-\hat{\Psi}_{2}\|_{F}}{2}\leq\frac{\|\hat{\Psi}_{1}-\hat{\Psi}\|_{F}}{2}+\epsilon_{2},\qquad\forall\text{ CPTP maps }\Psi.$ (5.228)
@@
Even though $\hat{\Psi}_{2}$ has desirable distance properties, as it does not achieve exact convergence to the projection, it may not correspond to a CPTP map. However, as our final projection was onto the set $C_{2}$, $\hat{\Psi}_{2}$ is Hermitian and satisfies the affine constraints. Thus, the only remaining condition necessary to obtain a proper hypothesis is to obtain a positive semidefinite hypothesis. To ensure that our final hypothesis is PSD while still retaining the other properties, we take inspiration from [surawy2022projected] and mix it with the maximally mixed state. In particular, for a suitably chosen $p>0$, our output is
@@
where $I$ is the $D\times D$ identity operator. To understand the intuition behind mixing, we first note that the maximally mixed state can be viewed as the Fourier coefficient matrix of a CPTP map that applies a Pauli operator uniformly at random. Thus, it must satisfy all required constraints. Then, we note that if $p$ is taken large enough, $\hat{\Psi}_{3}$ also satisfies the constraints. This can be seen by noting that the convex combination of two points satisfying the affine constraints will also satisfy them. For the positive semidefiniteness, we note that $\hat{\Psi}_{2}$ is Hermitian, and we only need to take $p$ large enough to make the most negative eigenvalue of $\hat{\Psi}_{2}$ map to 0 in the eigenspectrum of $\hat{\Psi_{3}}$. Using the fact that $\hat{\Psi}_{2}$ is not too far from a PSD matrix (Equation 5.226), we will then show that $p$ does not need to be too large.
@@
We can now upper bound $|\lambda^{\star}|$ as follows. First, recall that projection of a Hermitian operator onto the positive semidefinite cone is well-known (see, e.g., [henrion2012projection]). In particular, this projection works by taking the positive component of the Hermitian operator’s eigenspectrum and setting the negative eigenvalues to 0. For a Hermitian $A$, we denote this by $(A)^{+}$. Then, we have
@@
$=\frac{p}{2}\mathinner{\!\left\lVert\hat{\Psi}_{2}-\frac{I}{D}\right\rVert}_{F}$ (5.238)
$\leq\frac{p}{2}\left(\|\hat{\Psi}_{2}\|_{F}+\mathinner{\!\left\lVert\frac{I}{D}\right\rVert}_{F}\right)$ (5.239)
$\leq\frac{p}{2}\left(\|\hat{\Psi}_{2}-\hat{\Psi}^{*}\|_{F}+\|\hat{\Psi}^{*}\|_{F}+\mathinner{\!\left\lVert\frac{I}{D}\right\rVert}_{F}\right)$ (5.240)